OpenStudy (czesc):

INSTANT FAN//MEDAL What is distance around the graph of r = 12 sin θ ?

OpenStudy (mathmale):

Could you possibly draw a diagram of this situation? Rather than "distance around the graph," do you mean "perimeter" or "arc length?"

OpenStudy (mathmale):

Have you calculated arc length before? If so, using which formula or formulas?

OpenStudy (astrophysics):

I'll give you a hint on how to draw the graph as it can be pretty tricky at times, start with \[\sin \theta = \frac{ y }{ r }\] then we have \[r=12 \left( \frac{ y }{ r } \right)\] and also note that \[r^2 = x^2+y^2\] see what you can do with this.

OpenStudy (czesc):

@mathmale I'm assuming they want circumference I don't know what that means to be quite honest, and no I have not calculated arc length before. Thanks for the response. @Astrophysics Thank you I'll start graphing.

OpenStudy (czesc):